• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:29:00 on October 7,2002

    Entry number 87566

    keyword=optics target position

    The current target position is optics target in beam position. Dave Meekins told me that he re-adjusted the position for optics target and it should be in the correct position now.